# Imperial College London

### v1

@@ VI Conclusion
Analog computing is undergoing a revival, promising low-power and massively parallel computation capabilities for applications in signal processing and communications. In this paper, we introduce the concept of MiLAC as a generic analog computer that linearly operates with microwave signals, and analyze its application to signal processing for communications. A MiLAC can generally be modeled as a multiport microwave network made of tunable impedance (or admittance) components. At the ports of such a microwave network, the input signals are applied and the output signals are read. By using rigorous microwave theory, we derive the expression of the output of a MiLAC as a function of its input and the values of its admittance components. We show that a MiLAC can efficiently compute the LMMSE estimator in the analog domain, with significantly reduced computational complexity over conventional digital computing. In particular, it can be performed with complexity growing with the square of the input size, rather than the cube as with digital computing. This remarkable reduction in computational complexity is expected to play a key role in enabling beamforming and detection in gigantic MIMO communications.
@@ VI-1 Scalar Product
VI-2 Matrix-Vector Product
@@
VI-3 Matrix-Matrix Product
@@
VI-4 Matrix Inversion
Given an invertible matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, the algorithm commonly used to compute the matrix inverse $\mathbf{A}^{-1}$ is the Gauss’s method [22, Chapter 1]. Thus, we regard the complexity of matrix inversion as equal to the complexity of the Gauss’s method algorithm. It can be shown following [22, Chapter 1] that the Gauss’s method requires $N(N+1)/2$ divisions, $(2N^{3}+3N^{2}-5N)/6$ multiplications, and $(2N^{3}+3N^{2}-5N)/6$ subtractions. Consequently, the algorithm’s complexity is dominated by approximately $N^{3}/3$ multiplications and $N^{3}/3$ subtractions, which require a total of $38N^{3}/$ 3 real operations.

### v2

@@ VII Conclusion
Analog computing is undergoing a revival, promising low-power and massively parallel computation capabilities for applications in signal processing and communications. In Part I of this paper, we introduce the concept of MiLAC as a generic analog computer that linearly operates with microwave signals, and analyze its application to signal processing. A MiLAC can generally be modeled as a multiport microwave network made of tunable impedance (or admittance) components. At the ports of such a microwave network, the input signals are applied and the output signals are read. By using rigorous microwave theory, we derive the expression of the output of a MiLAC as a function of its input and the values of its admittance components. We show that a MiLAC can efficiently compute the LMMSE estimator and invert an arbitrary matrix in the analog domain, with significantly reduced computational complexity over conventional digital computing. In particular, matrix inversion can be performed with complexity growing with the square of the matrix size, rather than the cube as with digital computing. The primitive operations computable with a MiLAC can serve as building blocks for implementing more complex algorithms, such as the Kalman filter, directly in the analog domain, achieving a substantial decrease in computational complexity. Finally, to facilitate the practical implementation of a MiLAC, we present a design based on lossless admittance components, avoiding losses and costly active RF components. In Part II of this paper, we discuss the application of MiLAC to wireless communications, showing that it can enable gigantic MIMO communications by performing computations and beamforming in the analog domain.
@@ VII-1 Scalar Product
VII-2 Matrix-Vector Product
@@
VII-3 Matrix-Matrix Product
@@
VII-4 Matrix Inversion
Given an invertible matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, the algorithm commonly used to compute the matrix inverse $\mathbf{A}^{-1}$ is the Gauss’s method [44, Chapter 1]. Thus, we regard the complexity of matrix inversion as equal to the complexity of the Gauss’s method algorithm. It can be shown following [44, Chapter 1] that the Gauss’s method requires $N(N+1)/2$ divisions, $(2N^{3}+3N^{2}-5N)/6$ multiplications, and $(2N^{3}+3N^{2}-5N)/6$ subtractions. Consequently, the algorithm’s complexity is dominated by approximately $N^{3}/3$ multiplications and $N^{3}/3$ subtractions, which require a total of $38N^{3}/$ 3 real operations.